# Revisiting the question of what instantaneous normal modes tell us about liquid dynamics

### v1

@@ Abstract
The absence of a well-defined equilibrium reference configuration and the inevitable frequent atomic rearrangements have long obstructed the achievement of a complete atomic-level understanding of liquid dynamics and properties, a challenge that continues to be unresolved. The instantaneous normal mode (INM) approach, based on the diagonalization of the potential energy Hessian matrix in instantaneous liquid configurations, has been shown to be a promising starting point to predict thermodynamic and dynamical properties of liquids but presents several conceptual difficulties that remain to be addressed. More in general, due to the inability of capturing anharmonic effects, what INMs can tell us about liquid dynamics remains an open question. In this work, we provide a general set of “experimental facts” about the INM spectrum of liquids by performing a comprehensive INM analysis of several simulated systems, including Ar, Xe, N2, CS2, Ga and Pb, in a large range of temperatures from the solid to the gas phase. We first study the INM density of states (DOS) and compare it to the density of state function obtained from the velocity auto-correlation function. Secondly, we analyze the temperature dependence of the fraction of unstable modes and of the low-frequency slope of the INM DOS, in search of possible universal behaviors. We then explore the connection between INMs and other dynamical and structural properties of liquids including the liquid-like to gas-like dynamical crossover and the momentum gap of collective shear waves. Moreover, we investigate the INM spectrum at low temperature and its properties upon approaching the solid phase, revealing the existence of a large fraction of unstable modes also in crystalline solids. Finally, we verify the existence of a recently discussed cusp-like singularity in the INM eigenvalue spectrum and reveal its complex behavior upon dialing temperature that challenges the existing theoretical models.
@@ I Introduction
Normal modes describe the coherent motion of atoms oscillating at the same frequency and they represent the basis for the study of vibrations in many physical systems from buildings and bridges to molecules and solid state materials Cui and Bahar (2005); Dykeman and Sankey (2010). The most general dynamics in the linear regime can be indeed reconstructed using a superposition of normal modes that act as building blocks for thermodynamic and transport properties as well. A paradigmatic example is that of harmonic crystals in which the normal mode analysis provides a complete atomic-scale description Moon (2023), and leads to the well-established and successful Debye Debye (1912) and Einstein Einstein (1907) models.
@@
An analogous microscopic description of the physical properties of liquids incurs in several difficulties (see Moon for an extensive discussion) and remains unachieved Cabrillo et al. (2002). Consistent with Maxwell’s idea Maxwell (1867) that liquids behave as solids below a certain time-scale (Maxwell relaxation time), in 1967 Zwanzig proposed that normal modes could also exist in liquids, and be a fundamental part of the dynamics Zwanzig (1967). This hypothesis was quickly verified by Rahman, Mandell, and McTague Rahman et al. (1976) using a LJ simulated model, giving birth to the concept of “instantaneous normal modes” (INM), that was developed later on to great extent by many Authors Keyes (1997a).
@@
INM can be obtained by probing the potential energy surface (PES) at any instant of time and diagonalizing the corresponding instantaneous Hessian matrix that encodes all the information about the local curvature of the PES. In this way, the INM density of states (DOS) can be defined by averaging on several of these instantaneous configurations. Despite nowadays this procedure can be easily achieved using computer simulations, what INM are and are not remains a challenging question Stratt (1995). In particular, whether the harmonic oscillator picture is just a mere mathematical formal analogy or some kind of coherent oscillatory motion occurs in liquids is a topic of debate. The INM analysis assumes that interactions are harmonic (fully captured by the second derivative of the potential – Hessian matrix), which is a very questionable assumption in liquids. Nevertheless, this crude approximation might have some validity if the Maxwell picture, where short time dynamics is elastic in nature, is accepted. Importantly, short time dynamics has repeatedly been found to be directly linked to long time dynamics so that the limited applicability of the INM analysis might conceivably be extended to longer time relaxation processes. [^1] We would like to thank Jack F. Douglas for suggesting this point to us.
@@
Among the various difficulties in this program, it was early realized by Rahman Mandell, and McTague Rahman et al. (1976) that, despite the INM density of states displays striking similarities with the Fourier transform of the velocity auto-correlation function (VACF), the INM procedure inevitably leads to the presence of negative eigenvalues corresponding to regions in the PES with locally negative curvature. These negative eigenvalues translate into imaginary frequencies, i.e., “unstable modes”. INM can be therefore divided into stable (positive eigenvalues and real frequencies) and unstable (negative eigenvalues and imaginary frequencies), and the same separation can be applied to the corresponding INM DOS. Let us immediately remark that the “density of states” of unstable INM cannot be directly interpreted as a vibrational density of states. More in general, unstable modes cannot be identified with collective atomic vibrations and their physical interpretation appears more subtle and related to the relaxational processes, active in liquids but suppressed in crystals.
@@
The fact that liquids display solid-like oscillatory dynamics at large frequencies, or equivalently short times, it is not a mystery nor a surprise Vallauri and Bermejo (1995), and it is experimentally verified, e.g., Giordano and Monaco (2010). These fast dynamics can be well captured by the INM analysis and by the concept of stable INM, whose density of states is in good agreement with the Fourier transform of the VACF above a certain cutoff frequency (see for example Seeley et al. (1992)). On the other hand, it is well-know that these two quantities are different at low frequencies (compared to an inverse characteristic time $1/\tau$). As a concrete example, the zero frequency value of the INM DOS vanishes while the corresponding limit in the VACF gives the non-zero self-diffusion constant. This is just a manifestation that a harmonic description of liquids cannot apply at long times because of the inevitably anharmonic nature of the potential landscape David and Stratt (1998) (some anharmonic extensions of the INM approach have been proposed in the past, e.g., Krämer et al. (1998)). It was therefore a surprising discovery Keyes (1994a, 1995) that INM, and in particular unstable ones, can still carry important information about the late-time liquid dynamics and that the self-diffusion constant can be reconstructed from the distribution of some of them (even if a concrete characterization of the adjective “some” is a challenging task, see Gezelter et al. (1997) and related discussions Li et al. (1998); Chowdhary and Keyes (2002); Gezelter et al. (1999); Clapa et al. (2012a); Madan et al. (1990a)). More recently, it has been proposed that another transport property of liquids – the shear viscosity – can be reconstructed from unstable INMs Huang et al. (2024).
In frequency domain, the INM density of states exhibits interesting features that have been subject of intense investigation in the past decades. Particular attention has been devoted to understanding the frequency and temperature dependence of the unstable INM DOS Keyes et al. (1997); Vijayadamodar and Nitzan (1995); Zürcher and Keyes (1997); Xu et al. (2023), which is commonly represented on the negative frequency axes upon performing a variable transformation $\omega=-\sqrt{|\lambda|}$ with $\lambda$ the (negative) eigenvalues. Among the various features, the unstable INM DOS displays a universal linear in frequency scaling at low frequency that is observed in uncountable simulations and derived within several theoretical frameworks Xu and Stratt (1989); Wu and Loring (1993); Wan and Stratt (1994); Zaccone and Baggioli (2021); Schirmacher et al. (2022); Li and Keyes (1999); Zürcher and Keyes (1997); Wu and Loring (1992). It is nowadays clear that the emergence of this scaling is intimately connected to the existence of unstable modes. Moreover, by construction the stable and unstable INM DOS are symmetric upon the transformation $\omega\rightarrow-\omega$, implying that also the stable branch of the INM DOS displays a linear scaling with the same temperature dependent coefficient. Additional interesting properties emerge when the INM DOS is represented in function of the eigenvalues, rather than frequencies. In particular, it was early observed Sastry et al. (2001); Taraskin and Elliott (2002), and recently re-considered Schirmacher et al. (2022); Mossa et al. (2023), that the eigenvalue distribution presents a cusp-like singularity at $\lambda=0$ that remains hidden in the frequency representation but carries important information about liquid dynamics as well.
@@
The relation (if any) between the INM DOS, the Fourier transform of the VACF (sometimes called the “density of state function”) and the experimentally measurable density of states of liquids remains an important open question. The density of states of liquids at low energies is much less explored experimentally compared to the solid counterpart as the standard experimental probes incur in several difficulties. Until very recently, only few works explored the low-frequency DOS of liquids using inelastic neutron scattering techniques, i.e., Ref. Phillips et al. (1989) for liquid selenium and Ref. Dawidowski et al. (2000) for heavy water. More recently, the low-frequency experimental DOS of different liquid systems was investigated in Ref. Stamper et al. (2022) and a universal linear scaling was observed in agreement with the expectations from the INM analysis. Furthermore, the temperature dependence of the linear scaling of the experimental DOS was studied in Ref. Jin et al. (2024) and compared to the linear scaling of the INM DOS. Both scaling coefficients were found to follow the same exponential dependence on the inverse temperature; nevertheless, the corresponding activation energies were found to be different. Despite these encouraging developments, how to define “a” liquid density of states, and if such a quantity ultimately exists, remain works in progress.
@@
Arguably, the difficulties in performing a normal mode analysis for liquids, and in identifying a well-defined and unique density of states, are reflected in the challenge of achieving a microscopic description of liquid thermodynamics and collective dynamics Trachenko and Brazhkin (2015a). In the last years, several works Baggioli and Zaccone (2021); Xu et al. (2023); Moon et al. (2024, 2023) have attempted to use INMs to describe the thermodynamic properties of liquids and in particular their heat capacity. From these analyses, it has emerged that unstable modes cannot be ignored in liquids as considering only stable normal modes would lead to erroneous conclusions. This is somehow consistent with the older two-phase description of liquids, e.g. Lin et al. (2003), when the gas-like modes are related to the unstable INM modes as done in Ref. Moon et al. (2024). Interestingly, unstable INM have started to play an important role in glasses and amorphous solids as well Kriuchevskyi et al. (2022); Oyama et al. (2021). For example, unstable INM can be used to locate the onset of glass transition Bembenek and Laird (1995). Finally, it is not yet clear how the INM picture developed for liquids merges into the more traditional normal mode description of solids, for which Debye’s law has to be recovered. It is indeed known Pallikara et al. (2022), but often discarded, that also crystalline systems at finite temperature possess a non negligible number of unstable modes.
@@
In summary, there is increasing evidence that instantaneous normal modes might provide a basis for a microscopic description of liquid properties from their atomic-level motion, to their thermodynamics and collective properties. Nevertheless, the question of what we can learn from INM (in particular the unstable ones) is still open, see e.g. Zhang et al. (2019). In this work, we revisit this question and we provide a comprehensive INM analysis for many liquid systems in order to collect a series of “experimental facts” (hopefully) useful to build the long-sought microscopic theory of liquids.
@@ II.1 Interaction potential
In this work, all simulations are conducted via the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) Plimpton (1995). The computer simulation of metals was conducted using the pair interparticle interaction method, along with the incorporation of a more advanced embedded atom model (EAM) potential technique. In the EAM, an additional form of collective interaction is introduced alongside the conventional pair interparticle potentials. The expression for the potential energy of a metal is given as
@@
where $\Psi$$(r_{ij})$ is the contribution to electron density of neighbor atom $j$. The second term in Eq. (1) represents the usual pair potential. In this work, the interactions between metal crystals of Ga Belashchenko (2012) and Pd Wang et al. (2018) atoms are described with EAM potentials independently, which are widely applied in respective metallic systems Mokshin et al. (2015); Chtchelkatchev et al. (2016); Belashchenko (2017); Wonglakhon et al. (2021). The rest of non-metal systems of particles interact via the 12-6 Lennard-Jones (LJ) potential,
@@
where $K_{NN}$ is the force constant which is measure of interatomic bond between nitrogen atoms in N2 molecule, $r$ is the bond length, and $r_{e}$ is the bond length at equilibrium. We used a bond spring constant $K_{NN}$ = 13800 KJ/(mol) and a bond length $r_{e}=1.0975$. At the same time, the harmonic type potential for CS2 molecules S.-B. Zhu and Robinson (1988) is given by
@@
Ar Mourits and Rummens (1977) ArAr 3.465 0.225 8
Xe Mourits and Rummens (1977) XeXe 4.05 0.457 10
N2 Gao and Weiner (1990) NN 3.31 0.0715 12
CS2 Tildesley and Madden (1981) CC 3.35 0.1017

### v2

@@ Abstract
The lack of a well-defined equilibrium reference configuration has long hindered a comprehensive atomic-level understanding of liquid dynamics and properties. The Instantaneous Normal Mode (INM) approach, which involves diagonalizing the Hessian matrix of potential energy in instantaneous liquid configurations, has emerged as a promising framework in this direction. However, several conceptual challenges remain, particularly related to the approach’s inability to capture anharmonic effects. In this study, we present a set of "experimental facts" through a comprehensive INM analysis of simulated systems, including Ar, Xe, N2, CS2, Ga, and Pb, across a wide temperature range from the solid to gas phase. First, we examine the INM density of states (DOS) and compare it to the DOS obtained from the velocity auto-correlation function. We then analyze the temperature dependence of the fraction of unstable modes and the low-frequency slope of the INM DOS in search of potential universal behaviors. Furthermore, we explore the relationship between INMs and other properties of liquids, including the liquid-like to gas-like dynamical crossover and the momentum gap of collective shear waves. In addition, we investigate the INM spectrum at low temperatures as the system approaches the solid phase, revealing a significant fraction of unstable modes even in crystalline solids. Finally, we confirm the existence of a recently discussed cusp-like singularity in the INM eigenvalue spectrum and uncover its complex temperature-dependent behavior, challenging current theoretical models.
@@ I Introduction
Normal modes describe the coherent motion of atoms oscillating at the same frequency and they represent the basis for the study of vibrations in many physical systems from buildings and bridges to molecules and solid state materials [1, 2]. The most general dynamics in the linear regime can be indeed reconstructed using a superposition of normal modes that act as building blocks for thermodynamic and transport properties as well. A paradigmatic example is that of harmonic crystals in which the normal mode analysis provides a complete atomic-scale description [3], and leads to the well-established and successful Debye [4] and Einstein [5] models.
@@
An analogous microscopic description of the physical properties of liquids incurs in several difficulties (see [6] for an extensive discussion) and remains unachieved [7]. Based on molecular flow in liquids, Frenkel interpreted the Maxwell relaxation time [8] as a characteristic time between particle jumps and proposed that liquids are solid-like at shorter times (see [9] for a complete description of these historical developments). Importantly, Frenkel was also the first to discuss normal modes and their different propagation regimes in liquids (see also important contributions by Zwanzig [10]). The hypothesis that normal modes could also exist in liquids, and be a fundamental part of the dynamics, was quickly verified by Rahman, Mandell, and McTague [11] using a LJ simulated model, giving birth to the concept of “instantaneous normal modes” (INMs), that was developed later on to great extent by many authors.[12].
@@
INMs can be obtained by probing the potential energy surface (PES) at any instant of time and diagonalizing the corresponding instantaneous Hessian matrix that encodes all the information about the local curvature of the PES. In this way, the INM density of states (DOS) can be defined by averaging on several of these instantaneous configurations. Although, at present, this procedure can be easily achieved using computer simulations, what INM are and are not remains a challenging question [13]. In particular, whether the harmonic oscillator picture is just a mere mathematical formal analogy or some kind of coherent oscillatory motion occurs in liquids is a topic of debate. The INM analysis assumes that interactions are harmonic (fully captured by the second derivative of the potential – Hessian matrix), which is a very questionable assumption in liquids. Nevertheless, this crude approximation might have some validity if the Maxwell picture, where short time dynamics is elastic in nature, is accepted. Importantly, short time dynamics has repeatedly been found to be directly linked to long time dynamics so that the limited applicability of the INM analysis might conceivably be extended to longer time relaxation processes.
@@
Among the various difficulties in this program, it was early realized by Rahman Mandell, and McTague [11] that, although the INM density of states displays striking similarities with the Fourier transform of the velocity auto-correlation function (VACF), the INM procedure inevitably leads to the presence of negative eigenvalues corresponding to regions in the PES with locally negative curvature. These negative eigenvalues translate into imaginary frequencies, i.e., “unstable modes”. INMs can be, therefore, divided into stable (positive eigenvalues and real frequencies) and unstable (negative eigenvalues and imaginary frequencies), and the same separation can be applied to the corresponding INM DOS. Let us immediately remark that the “density of states” of unstable INM cannot be directly interpreted as a vibrational density of states. More in general, unstable modes cannot be identified with collective atomic vibrations and their physical interpretation appears more subtle and related to the relaxational processes, active in liquids but suppressed in crystals. It is interesting to notice that thermal fluctuations, moving atoms away from equilibrium, can directly induce the emergence of imaginary modes in the harmonic approximation. As a direct example, let us consider a LJ potential $V=A/r^{12}-B/r^{6}$: the equilibrium position is $r^{*}=2^{1/6}(A/B)^{1/6}$ and the second derivative is negative for $r$ larger than about $10\%$ of $r^{*}$. This gives an imaginary frequency for a finite fraction of atoms that are not performing any diffusing dynamics but rather oscillating around their equilibrium position due to thermal motion.
@@
More in general, the fact that liquids display solid-like oscillatory dynamics at large frequencies, or equivalently short times, it is not a mystery nor a surprise [14], and it is experimentally verified, e.g., [15]. These fast dynamics can be well captured by the INM analysis and by the concept of stable INMs, whose density of states is in good agreement with the Fourier transform of the VACF above a certain cutoff frequency (see, for example [16]). Meanwhile, it is well-know that these two quantities are different at low frequencies (compared to an inverse characteristic time $1/\tau$). As a concrete example, the zero frequency value of the INM DOS vanishes while the corresponding limit in the VACF gives the non-zero self-diffusion constant. This is just a manifestation that a harmonic description of liquids cannot apply at long times because of the inevitably anharmonic nature of the potential landscape [17] (some anharmonic extensions of the INM approach have been proposed in the past, e.g., [18]). It was therefore a surprising discovery [19, 20] that INMs, and in particular unstable ones, can still carry important information about the late-time liquid dynamics and that the self-diffusion constant can be reconstructed from the distribution of some of them (even if a concrete characterization of the adjective “some” is a challenging task, see [21] and related discussions [22, 23, 24, 25, 26]). More recently, it has been proposed that another transport property of liquids – the shear viscosity – can be reconstructed from unstable INMs [27].
In frequency domain, the INM density of states exhibits interesting features that have been the subject of intense investigation over the past decades. Particular attention has been devoted to understanding the frequency and temperature dependence of the unstable INM DOS [28, 29, 30, 31], which is commonly represented on the negative frequency axes upon performing a variable transformation $\omega=-\sqrt{|\lambda|}$ with $\lambda$ being the (negative) eigenvalues. Among the various features, the unstable INM DOS displays a universal linear in frequency scaling at low frequencies that is observed in uncountable simulations and derived within several theoretical frameworks [32, 33, 34, 35, 36, 37, 30, 38]. At present, it is clear that the emergence of this scaling is intimately connected to the existence of unstable modes. In order to avoid any misunderstanding, we stress that the linear scaling of the INM DOS is not in contradiction with numerous experimental results confirming the presence of linearly dispersing acoustic hydrodynamic modes in liquids (e.g., [39, 40]) that, on the other hand, correspond to a Debye quadratic DOS. In fact, the linear term in the DOS does not arise from propagating modes but rather from overdamped, and quasi-localized, excitations that are abundant in liquids as a consequence of many relaxational mechanisms (see, for example, [41] for a recent discussion on this point). The quadratic DOS is recovered at low temperatures in the solid state, but only if non-diffusive unstable modes are filtered out (see, for example, [42] for a particular method to do that).
Additional interesting properties emerge when the INM DOS is represented as a function of the eigenvalues, rather than frequencies. In particular, it was early observed [43, 44], and recently re-considered [36, 45], that the eigenvalue distribution presents a cusp-like singularity at $\lambda=0$ that remains hidden in the frequency representation but carries important information about liquid dynamics as well. We also notice that the properties of INM are strongly related to the topology of the energy landscape. In particular, unstable modes are controlled by the statistics of the saddle points in the potential energy landscape (PEL). The role played by saddles in the supercooled liquid dynamics has been explored in detail in the past (e.g., [46, 47]). These analyses have revealed that the order of the saddles decreases with the temperature and vanishes at the mode coupling temperature $T_{MCT}$, below which the dynamics are mainly localized within the basins. At the same time, the fraction of delocalized unstable modes vanishes at the same temperature that, thus, corresponds to a localization transition for the unstable INM [25, 48].
@@
The relation (if any) between the INM DOS, the Fourier transform of the VACF (sometimes called the “density of state function”) and the experimentally measurable density of states of liquids remains an important open question. We notice that a connection between normal modes in liquids and the VACF, and consequently the diffusion constant and the viscosity, was already advanced by Zwanzig in [49]. The density of states of liquids at low energies is much less explored experimentally compared to the solid counterpart as standard experimental probes incur in several difficulties. Until very recently, only few studies explored the low-frequency DOS of liquids using inelastic neutron scattering techniques, i.e., Ref. [50] for liquid selenium and Ref. [51] for heavy water. More recently, the low-frequency experimental DOS of different liquid systems was investigated in Ref. [52] and a universal linear scaling was observed in agreement with the expectations from the INM analysis. Furthermore, the temperature dependence of the linear scaling of the experimental DOS was studied in Ref. [53] and compared to the linear scaling of the INM DOS. Both scaling coefficients were found to follow the same exponential dependence on the inverse temperature; nevertheless, the corresponding activation energies were found to be different. Very recently, a liquid-like linear in frequency DOS has been experimentally observed in a solid-state lithium electrolyte at high temperatures [54]. Moreover, a crossover between a Debye DOS and a linear DOS has been obtained in simulations of amorphous hafnia upon crossing the glass transition temperature [55]. Despite these encouraging developments, how to define “a” liquid density of states, and if such a quantity ultimately exists, remains work in progress.
@@
Arguably, the difficulties in performing a normal mode analysis for liquids, and in identifying a well-defined and unique density of states, are reflected in the challenge of achieving a microscopic description of liquid thermodynamics and collective dynamics [56]. over the past years, several studies [57, 31, 58, 59] have been attempted to use INMs to describe the thermodynamic properties of liquids and in particular their heat capacity. From these analyses, it has emerged that unstable modes cannot be ignored in liquids as considering only stable normal modes would lead to erroneous conclusions. This is somehow consistent with the older two-phase description of liquids, e.g. [60], when the gas-like modes are related to the unstable INM modes as done in Ref. [58]. Interestingly, unstable INM have started to play an important role in glasses and amorphous solids as well [61, 62]. For example, unstable INM have been advocated in the past to locate the onset of glass transition [63]. Finally, it is not yet clear how the INM picture developed for liquids merges into the more traditional normal mode description of solids, for which Debye’s law has to be recovered. It is indeed known [64], but often discarded, that also crystalline systems at finite temperature possess a non negligible number of unstable modes. It is not completely understood whether these remaining modes are spurious or contain valuable physical information. In fact, one possible interpretation is that these modes are related to local anharmonicities that do not correspond to crossing barriers and moving to a different basin (hence, they do not contribute to diffusion). In other words, these are not necessarily spurious modes but rather indicators of anharmonicity within a potential well. On the other hand, several methods have been developed to filter them out, as for example considering only pure translation modes [42]. A final verdict on the nature of these modes is still an open question that goes beyond the scope of this work.
@@
In summary, there is increasing evidence that instantaneous normal modes might provide a basis for a microscopic description of liquid properties from their atomic-level motion, to their thermodynamics and collective properties. Nevertheless, the question of what we can learn from INM (in particular the unstable ones) is still open, see e.g. [65]. In this work, we revisit this question and we provide a comprehensive INM analysis for many liquid systems in order to collect a series of “experimental facts” (hopefully) useful to build the long-sought microscopic theory of liquids.
@@ II.1 Interaction potential
In this work, all simulations are conducted via the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [66]. The computer simulation of metals was conducted using the pair interparticle interaction method, along with the incorporation of a more advanced embedded atom model (EAM) potential technique. In the EAM, an additional form of collective interaction is introduced alongside the conventional pair interparticle potentials. The expression for the potential energy of a metal is given as
@@
where $\Psi$$(r_{ij})$ is the contribution to electron density of neighbor atom $j$. The second term in Eq. (1) represents the usual pair potential. In this work, the interactions between metal crystals of Ga [67] and Pd [68] atoms are described with EAM potentials independently, which are widely applied in respective metallic systems [69, 70, 71, 72]. The rest of non-metal systems of particles interact via the 12-6 Lennard-Jones (LJ) potential,
@@
where $K_{NN}$ is the force constant which is measure of interatomic bond between nitrogen atoms in N2 molecule, $r$ is the bond length, and $r_{e}$ is the bond length at equilibrium. We used a bond spring constant $K_{NN}$ = 13800 KJ/(mol) and a bond length $r_{e}=1.0975$. At the same time, the harmonic type potential for CS2 molecules [73] is given by
@@
Ar [74] ArAr 3.465 0.225 8
Xe [74] XeXe 4.05 0.457 10
N2 [75] NN 3.31 0.0715 12
CS2 [76] CC 3.35 0.1017